# Science Opportunities of Wet Extreme Mass-Ratio Inspirals

### v2

@@ I Introduction
The structure of this work is organized as follows. In Sec. II.1, we discuss the formation and population of wet EMRIs, including their formation mechanisms in accretion disks and population estimates based on updated SMBH mass functions. In Sec. II.2, we explore transient electromagnetic (EM) counterparts, focusing on the connection between wet EMRIs and QPEs. In Sec. II.3, we address how wet EMRIs can be used to calibrate other methods of probing BH mass and spin. In Sec. II.4, we examine the potential of wet EMRIs to test accretion disk and jet models. Finally, in Sec. II.5, we discuss the application of wet EMRIs as cosmic sirens. Resolvable host AGNs enable their use as bright sirens (yielding luminosity distance and redshift), while unresolvable hosts result in dark sirens. Finally, in Sec. III, we possible uncertainties and caveats in this analysis, and motivate further studies in this emerging field.
@@ II.1 Population estimation
$\alpha$-disk $10^{6}$ 2700 120 30
$^{7} 1100$ 40 9
$10^{8}$ 390 13 3
$\beta$-disk $10^{6}$ 3500 150 32
$10^{7}$ 1200 50 10
$10^{8}$ 390 13 3
@@
    $\frac{dN_{\bullet}}{d\log M_{\bullet}}=0.01\left(\frac{M_{\bullet}}{3\times 10^{6}M_{\odot}}\right)^{-0.3}\,\text{Mpc}^{-3},$
@@
In both cases, $dN_{\bullet}/d\log M_{\bullet}$ represents the number density of SMBHs per logarithmic mass interval. These mass functions were used in our previous studies [22, 31] to estimate the abundance of SMBHs in the range $10^{4}\sim 10^{7}M_{\odot}$, which is particularly relevant for the formation and detection of EMRI. Our results demonstrated that wet EMRI could significantly enhance total and detectable EMRI rates, often dominating over dry EMRI.
@@ II.1.1 Population and detectability
To evaluate LISA detection rates and conduct the Fisher matrix analyses presented in this work, we adopt the Augmented Analytic Kludge (AAK) waveform model [54, 55, 56, 57, 58, 59], as implemented in Fast EMRI Waveforms (FEW) [59, 60, 61, 62, 63]. All analyses are carried out in a 14-dimensional parameter space that includes the mass and spin of the SMBH, the mass of the sBH, the luminosity distance, the initial semi-latus rectum, initial eccentricity, sky location, orbital inclination, and the initial orbital phases in $\phi,\theta,r$. For both dry and wet EMRIs, we assume isotropic distributions for sky location and SMBH spin direction. The three phase angles are sampled uniformly from $0,2\pi]$, coalescence times are randomly drawn from the interval $0,2\pi]$ years, and the cosine of the inclination angle is sampled uniformly from [1,1][-1,1], The initial eccentricity $e_{0}$ is assigned based on the formation channel: for dry EMRIs, $e_{0}$ is drawn from a uniform distribution in $[0,0.2]$, reflecting their dynamical origins. For wet EMRIs, we assume circular orbits with $e_{0}=0$, consistent with strong eccentricity damping in AGN disks.
@@
$\langle a,b\rangle=4\,\text{Re}\int_{f_{\text{min}}}^{f_{\text{max}}}\frac{\tilde{a}(f)\tilde{b}^{*}(f)}{P_{n}(f)}\,df.$ (7)
@@ II.1.2 Identifications of wet EMRIs and host AGNs
In wet environments, based on standard migration theory, strong interactions between sBHs and the accretion disks of rapidly accreting SMBHs—primarily via dynamical friction and disk-induced density waves—efficiently damp the vertical momentum and eccentricity of a sBH, aligning it with the disk plane and circularizing its orbit. Consequently, wet EMRIs typically exhibit near-zero eccentricities and low inclinations by the time they enter the LISA band.
@@
To evaluate the potential for identifying the AGN host of an EMRI event, we utilize Fisher matrix analysis, a well-established statistical method to estimate the precision of parameter measurements from observed signals. The Fisher information matrix, $\Gamma_{ij}$, is defined as
@@
where $h$ is the signal, $\lambda_{i}$ are the source parameters, and the inner product is defined by Eqn. 7. The covariance matrix, $\Sigma_{ij}=(\Gamma^{-1})_{ij}$, is obtained by inverting the Fisher matrix, with the diagonal elements, $\sigma_{\lambda_{i}}^{2}=\Sigma_{ii}$, represent the variances of the individual parameters, thereby quantifying the uncertainties and correlations associated with the source localization, including angular resolution.
@@
The solid angle uncertainty, $\Delta\Omega_{s}$, is derived from the angular components of the Fisher matrix, associated with the angular parameters $(\theta_{s},\phi_{s})$, and is given by (see also [71, 72])
$\Delta\Omega_{s}=2\pi\sin{\theta_{s}}\sqrt{\Sigma_{\theta_{s}\theta_{s}}\Sigma_{\phi_{s}\phi_{s}}-(\Sigma_{\theta_{s}\phi_{s}})^{2}}\,$, (9)
Combining this with the uncertainty in luminosity distance $d_{L}$, the three-dimensional sky localization volume is calculated as
$V_{\text{sky}}=\frac{1}{2}r^{3}(z)\,\sigma(\ln{d_{L}})\,\Delta\Omega_{s}\,$, (10)
where $r(z)$ is the comoving distance to the source at redshift $z$, and $\sigma(\ln{d_{L}})$ is the relative uncertainty in the luminosity distance measurement. For EMRIs detected by LISA, Fisher analysis indicates that $V_{\text{sky}}$ can be restricted to $\mathcal{O}(10^{2})\,\rm Mpc^{3}$ at moderate redshifts ($z<0.3$ [71]), depending on the SNR and the inclusion of higher order waveform harmonics, demonstrating the high localization precision of LISA.
To assess whether an EMRI can be classified as a bright siren, we adopt an average number of galaxies per comoving volume $\bar{n}_{\rm gal}=10^{-2}\rm Mpc^{-3}$ [73] and assume a conservative AGN fraction of $f_{\rm AGN}=1\%$. Although observational studies suggest that the AGN fraction is approximately $1\%$ at low redshift and may increases to $1-10\%$ at higher redshifts, we adopt a constant, redshift-independent AGN fraction of $1\%$ throughout this work to ensure a conservative and robust estimate.
@@ II.2 Transient Electromagnetic Counterparts: type II QPEs
where $\alpha$ is the standard viscosity coefficient in thin disks, $a$ is the dimensionless spin of the SMBH, $\dot{M}_{\bullet}$ is the SMBH accretion rate and the Eddington accretion rate $\dot{M}_{\bullet,\rm Edd}=M_{\bullet}/(5\times 10^{7}{\rm yr})$, and $h$ is the disk aspect ratio. If the accretion of the SMBH is coherent in a period much longer than $t_{\rm align}$, we expect an equatorial accretion disk and an equatorial EMRI when the EMRI enters the LISA band; therefore, no detectable EM transient is produced. Otherwise, if the gas feeding direction is random in each accretion episode with an accretion time scale shorter than $t_{\rm align}$, EM transients arising from EMRI and disk collisions are possible: for a sBH embedded in the disk, the Lense-Thirring precession tends to drive its orbit out of the disk plane, while the disk interaction tends to keep the sBH within the plane. We denote the critical radius as $r_{\rm dec}$ (as illustrated in FIG. 1), where the alignment timescale for the sBH orbit with the disk is equal to the Lense-Thirring precession timescale of the sBH. At smaller radii $r<r_{\rm dec}$, the sBH orbiting around the SMBH may hit the accretion disk and produce (quasi-)periodic EM transients, similar to some of the discussion of X-ray quasi-periodic eruptions (QPEs) [33, 34, 35, 36, 37, 38, 39, 40, 41, 42], where eruptions are highly likely due to collisions between a low-frequency ($f_{\rm obt}\sim 10^{-5}$ Hz) EMRI and an accretion disk formed in a recent TDE [80, 81, 82, 83, 84, 42, 41, 85, 86, 87, 88, 89, 90]. For clarity, we will call the TDE-associated QPEs as type I QPEs and the quasi-periodic EM transients produced by an EMRI crossing an AGN disk as type II QPEs:
@@
Notice that Swift J023017 is likely a type II QPE source as analyzed in [86], which is associated with a low luminosity AGN.
@@
$\frac{d\vec{L}}{dt}=\left(\vec{\omega}_{\rm LT}+\vec{\omega}_{\rm DI}\right)\times\vec{L}\ ,$ (14)
@@
$\omega_{\rm DI}:=-\frac{d\iota_{\rm sd}}{dt}={\rm min.}\left\{0.544\iota_{\rm sd},\frac{1.46h^{4}}{\sin^{3}(\iota_{\rm sd}/2)}\right\}\frac{1}{t_{\rm wav}}$ (15)
@@
$t_{\rm wav}=\frac{M}{\mu}\frac{M}{\Sigma\,r^{2}}\frac{h^{4}}{\omega_{\rm K}}\ ,$ (16)
@@
for $\beta$ disks, where we have defined $\mu_{30}=\mu/30M_{\odot},M_{\bullet,6}=M_{\bullet}/10^{6}M_{\odot},\dot{M}_{\bullet,0.1}=\dot{M}_{\bullet}/0.1\dot{M}_{\bullet,\rm Edd}$ and we have used analytic models of $\alpha$ and $\beta$ disks [13, 14]. The two numerical factors are nearly equal because the two disk models are similar at large radii where the gas pressure dominates the total pressure in the accretion disk. Due to the sharp $r$ dependence of the two rates $\omega_{\rm LT}$ and $\omega_{\rm DI}$, the decoupling radius $r_{\rm dec}$ turns out to be insensitive to $M_{\bullet},\alpha,\mu$, and only has a mild dependence on the SMBH accretion rate $\dot{M}_{\bullet}$. From Eqn. (14), it is straightforward to see that the sBH is nearly aligned with the disk ($\vec{L}/L\approx\vec{n}_{\rm disk}$) for $r>r_{\rm dec}$ and starts to precess around the SMBH spin axis at $r<r_{\rm dec}$ (as illustrated in FIG. 1). The transition between the two regimes is very sharp due to the sharp $r$ dependence of the two rates $\omega_{\rm LT}$ and $\omega_{\rm DI}$.
@@
respectively. Therefore, wet EMRI with a warped disk is a natural mechanism to generate type II QPEs. If type II QPEs are confirmed in future observations, their existence provides indirect evidence for warped disks and the chaotic accretion mode in AGNs.
As the EMRI enters the LISA sensitivity band with orbital radius $r\sim 10M_{\bullet}$ and orbital frequency $f_{\rm obt}\sim$ mHz, the energy budgets per collision for two disk models are quite different. In the $\alpha$-disk model, it scales as $\delta E_{\rm sBH}\propto r^{5/2}\propto f_{\rm obt}^{-5/3}$, that is say, the mHz QPEs will be $\mathcal{O}(10^{3})$ times weaker than their low-frequency counterparts. Although the identification of mHz type II QPEs is difficult in a sea of much stronger AGN variability, the joint GW measurement of the orbits may help identify the detection of mHz type II QPEs. If the accretion disk can be modeled as a $\beta$-disk, the energy budget per collision scales as $\delta E_{\rm sBH}\propto r^{2/5}\propto f_{\rm obt}^{-4/15}$, i.e. the mHz QPEs will be a few times weaker than their low-frequency counterparts. In this case, the identification of mHz QPEs is possible even without the input of GW observations.
@@
The other possible phenomenon to observe is the strong lensing of flares processed by the central SMBH, similar to self-lensing flares from SMBH binaries [93]. In this case, the strong lensing of the EMRI-produced flares is probably more likely to happen because the sBH is much closer to its host. As shown in [94], the strong lensing probability of a flare (either an EM flare or GWs) emitted at radius $r$ is approximately
@@
Therefore the chance of strong lensing of a single flare is around $\mathcal{O}(1\%)$ level, and the chance will be largely enhanced considering multi-flares are produced as the EMRI orbiting around the SMBH and crossing the accretion disk at different locations. Unlike the case of SMBH binaries with one of them hosting a mini-accretion disk where the background light source is stable and the variations in the light curve mainly come from when the mini disk gets lensed by the foreground BH, the light curve of mHz QPEs varies from flare to flare even without lensing. A more detailed study is required to confirm whether the lensing effect can be identified from the QPE light curve.

### v3

@@ I Introduction
This work is organized as follows: in Sec. II.1, we discuss the formation and population of wet EMRIs, including their formation mechanisms in accretion disks and population estimates based on updated SMBH mass functions. In Sec. II.2, we explore transient electromagnetic (EM) counterparts, focusing on the connection between wet EMRIs and QPEs. In Sec. II.3, we address how wet EMRIs can be used to calibrate other methods of probing BH mass and spin. In Sec. II.4, we examine the potential of wet EMRIs to test accretion disk and jet models. Finally, in Sec. II.5, we discuss the application of wet EMRIs as cosmic sirens. Resolvable host AGNs enable their use as bright sirens (yielding luminosity distance and redshift), while unresolvable hosts result in dark sirens. Finally, in Sec. III, we possible uncertainties and caveats in this analysis, and motivate further studies in this emerging field.
All masses (e.g., $M_{\bullet}$, $\mu$) used throughout the manuscript, including those appearing in the Fisher matrix estimates, are consistently expressed in the source frame unless explicitly stated otherwise.
@@ II.1 Population estimation
$\alpha$-disk $10^{6}$ 2700 90 5
$^{7} 1100$ 22 1.3
$10^{8}$ 390 9 0.4
$\beta$-disk $10^{6}$ 3500 100 6
$10^{7}$ 1200 30 1
$10^{8}$ 390 8 0.6
@@
    $\frac{dN_{\bullet}}{d\log M_{\bullet}}=0.01\left(\frac{M_{\bullet}}{3\times 10^{6}M_{\odot}}\right)^{-0.3}\,\text{Mpc}^{-3}\ .$
@@
In both cases, $dN_{\bullet}/d\log M_{\bullet}$ represents the number density of SMBHs per logarithmic mass interval. These mass functions were used in our previous studies [105, 102] to estimate the abundance of SMBHs in the range $10^{4}\sim 10^{7}M_{\odot}$, which is particularly relevant for the formation and detection of EMRIs. Our results demonstrated that wet EMRIs could significantly enhance total and detectable EMRI rates, often dominating over dry EMRIs.
@@ II.1.1 Population and detectability
To evaluate LISA detection rates and conduct the Fisher matrix analyses presented in this work, we adopt the Augmented Analytic Kludge (AAK) waveform model [15, 30, 29, 126, 31, 64], as implemented in Fast EMRI Waveforms (FEW) [64, 32, 46, 125, 65]. All analyses are carried out in a 14-dimensional parameter space that includes the mass and spin of the SMBH, the mass of the sBH, the luminosity distance, the initial semi-latus rectum, initial eccentricity, sky location, orbital inclination, and the initial orbital phases in $\phi,\theta,r$. For both dry and wet EMRIs, we assume isotropic distributions for sky location and SMBH spin direction. The three phase angles are sampled uniformly from $0,2\pi]$, and coalescence times are randomly drawn from the interval $[0,2]$ years. For wet EMRIs, the cosine of the inclination angle is sampled uniformly from the range $[0,0.1]$, reflecting alignment with the AGN disk, while for dry EMRIs it is drawn from a uniform distribution over [1,1][-1,1], allowing for isotropic orientations. The initial eccentricity $e_{0}$ is also assigned based on the formation channel: for dry EMRIs, $e_{0}$ is drawn from a uniform distribution in $[0,0.2]$, consistent with their dynamical origin in dense stellar environments. In contrast, wet EMRIs are assumed to start on circular orbits with $e_{0}=0$, reflecting strong eccentricity damping due to AGN disk interactions. We adopt a uniform spin distribution in the range $[0,0.99]$ for dry EMRIs, which reduces the detection rate for the dry channel by approximately $30\%$ compared to the case with fixed spin $a=0.98$, while for wet EMRIs we assume a fixed spin of $a=0.98$ to reflect efficient spin alignment in AGN disks.
@@
$\langle a,b\rangle=4\,\text{Re}\int_{f_{\text{min}}}^{f_{\text{max}}}\frac{\tilde{a}(f)\tilde{b}^{*}(f)}{P_{n}(f)}\,df\,$, (7)
@@ II.1.2 Identifications of wet EMRIs and host AGNs
In wet environments, based on standard migration theory, strong interactions between sBHs and the accretion disks of rapidly accreting SMBHs—primarily via dynamical friction and disk-induced density waves—efficiently damp the vertical momentum and eccentricity of an sBH, aligning it with the disk plane and circularizing its orbit. Consequently, wet EMRIs typically exhibit near-zero eccentricities and low inclinations by the time they enter the LISA band.
@@
To evaluate the potential for identifying the AGN host of an EMRI event, we utilize Fisher matrix analysis, a well-established statistical method to estimate the precision of parameter measurements for expected signals. The Fisher information matrix, $\Gamma_{ij}$, is defined as
@@
where $h$ is the signal, $\lambda_{i}$ are the source parameters, and the inner product is defined by Eq. 7. The covariance matrix, $\Sigma_{ij}=(\Gamma^{-1})_{ij}$, is obtained by inverting the Fisher matrix, with the diagonal elements, $\sigma_{\lambda_{i}}^{2}=\Sigma_{ii}$, represent the variances of the individual parameters, thereby quantifying the uncertainties and correlations associated with the source localization, including angular resolution.
Figure 1: Histograms of $1\sigma$ parameter uncertainties from the Fisher analysis, evaluated for $\sim 2000$ EMRI events with $\mathrm{SNR}>20$. Shown are distributions for $\Delta M_{\bullet}/M_{\bullet}$, $\Delta\mu/\mu$, $\Delta d_{L}/d_{L}$, $\Delta\iota$, $\Delta e_{0}$, and $\Delta a_{0}$. Since masses $M_{\bullet}$ and $\mu$ are reported in the source frame, their relative uncertainties are larger by a factor of $1+z$ compared to the detector frame.
To illustrate the expected precision in intrinsic and extrinsic parameters, we performed a Fisher matrix analysis over a population of $\sim 2000$ wet EMRI events with $\mathrm{SNR}>20$. FIG. 1 presents histograms of the $1\sigma$ uncertainties for the central black hole mass $M_{\bullet}$, compact object mass $\mu$ (source frame), luminosity distance $d_{L}$, orbital inclination $\iota$, initial eccentricity $e_{0}$, and SMBH spin $a_{0}$. These distributions reflect the statistical spread of parameter precisions expected in the LISA EMRI population.
These results demonstrate that LISA will achieve high-precision measurements for both intrinsic and extrinsic EMRI parameters, consistent with prior studies (e.g., Ref. [12]). Such precise parameter recovery will be instrumental in distinguishing EMRI formation channels and identifying host galaxies [127] .
@@
The sky localization volume quantifies the three-dimensional uncertainty in the inferred source location, combining angular position and luminosity distance. In the Fisher formalism, the solid angle uncertainty $\Delta\Omega_{s}$ is derived from the covariance matrix $\Sigma=\Gamma^{-1}$ and is given by [104, 63]
$\Delta\Omega_{s}=2\pi\sin{\theta_{s}}\sqrt{\Sigma_{\theta_{s}\theta_{s}}\Sigma_{\phi_{s}\phi_{s}}-(\Sigma_{\theta_{s}\phi_{s}})^{2}}\,.$ (9)
The full $68\%$ confidence localization volume in comoving coordinates is computed by including the correlation between angular position and luminosity distance. For a trivariate Gaussian distribution, the $68\%$ probability region corresponds to an ellipsoid with volume
$V_{\text{sky}}=28\,r^{3}(z)\,\sin{\theta_{s}}\,\sqrt{\det\left(\Sigma_{\ln d_{L},\,\theta_{s},\,\phi_{s}}\right)}\,$, (10)
where $r(z)$ is the comoving distance to redshift $z$, and $\Sigma_{\ln d_{L},\,\theta_{s},\,\phi_{s}}$ is the $3\times 3$ submatrix of the full covariance matrix corresponding to the parameters $(\ln d_{L},\theta_{s},\phi_{s})$. The prefactor 28 arises from the volume of a 3D Gaussian ellipsoid enclosing 68% of the total probability (see Appendix A for a detailed derivation). For EMRIs detected by LISA, Fisher analysis indicates that $V_{\text{sky}}$ can be restricted to $\mathcal{O}(10^{2})\,\rm Mpc^{3}$ at moderate redshifts ($z<0.3$ [104]), depending on the SNR and the inclusion of higher order waveform harmonics, demonstrating the high localization precision of LISA.
To assess whether an EMRI can be classified as a bright siren, we adopt an average number of galaxies per comoving volume $\bar{n}_{\rm gal}=10^{-2}\rm Mpc^{-3}$ [74] and assume a conservative AGN fraction of $f_{\rm AGN}=1\%$. Although observational studies suggest that the AGN fraction is approximately $1\%$ at low redshift and may increases to $1\%-10\%$ at higher redshifts, we adopt a constant, redshift-independent AGN fraction of $1\%$ throughout this work to ensure a conservative and robust estimate.
@@ II.2 Transient Electromagnetic Counterparts: type II QPEs
where $\alpha$ is the standard viscosity coefficient in thin disks, $a$ is the dimensionless spin of the SMBH, $\dot{M}_{\bullet}$ is the SMBH accretion rate and the Eddington accretion rate $\dot{M}_{\bullet,\rm Edd}=M_{\bullet}/(5\times 10^{7}{\rm yr})$, and $h$ is the disk aspect ratio. If the accretion of the SMBH is coherent in a period much longer than $t_{\rm align}$, we expect an equatorial accretion disk and an equatorial EMRI when the EMRI enters the LISA band; therefore, no detectable EM transient is produced. Otherwise, if the gas feeding direction is random in each accretion episode with an accretion time scale shorter than $t_{\rm align}$, EM transients arising from EMRI and disk collisions are possible: for an sBH embedded in the disk, the Lense-Thirring precession tends to drive its orbit out of the disk plane, while the disk interaction tends to keep the sBH within the plane. We denote the critical radius as $r_{\rm dec}$ (as illustrated in FIG. 2), where the alignment timescale for the sBH orbit with the disk is equal to the Lense-Thirring precession timescale of the sBH. At smaller radii $r<r_{\rm dec}$, the sBH orbiting around the SMBH may hit the accretion disk and produce (quasi-)periodic EM transients, similar to some of the discussion of X-ray quasi-periodic eruptions (QPEs) [128, 95, 53, 7, 8, 25, 43, 67, 58, 6], where eruptions are highly likely due to collisions between a low-frequency ($f_{\rm obt}\sim 10^{-5}$ Hz) EMRI and an accretion disk formed in a recent TDE [139, 78, 45, 129, 80, 6, 58, 147, 149, 148, 106, 24, 79]. For clarity, we will call the TDE-associated QPEs as type I QPEs and the quasi-periodic EM transients produced by an EMRI crossing an AGN disk as type II QPEs:
@@
Notice that Swift J023017 is likely a type II QPE source as analyzed in Ref. [149], which is associated with a low luminosity AGN.
@@
$\frac{d\vec{L}}{dt}=\left(\vec{\omega}_{\rm LT}+\vec{\omega}_{\rm DI}\right)\times\vec{L}\ .$ (14)
@@
$\omega_{\rm DI}:=-\frac{d\iota_{\rm sd}}{dt}={\rm min.}\left\{0.544\iota_{\rm sd},\frac{1.46h^{4}}{\sin^{3}(\iota_{\rm sd}/2)}\right\}\frac{1}{t_{\rm wav}}\,$, (15)
@@
$t_{\rm wav}=\frac{M}{\mu}\frac{M}{\Sigma\,r^{2}}\frac{h^{4}}{\omega_{\rm K}}\,.$ (16)
@@
for $\beta$ disks, where we have defined $\mu_{30}=\mu/30M_{\odot},M_{\bullet,6}=M_{\bullet}/10^{6}M_{\odot},\dot{M}_{\bullet,0.1}=\dot{M}_{\bullet}/0.1\dot{M}_{\bullet,\rm Edd}$ and we have used analytic models of $\alpha$ and $\beta$ disks [71, 143]. The two numerical factors are nearly equal because the two disk models are similar at large radii where the gas pressure dominates the total pressure in the accretion disk. Due to the sharp $r$ dependence of the two rates $\omega_{\rm LT}$ and $\omega_{\rm DI}$, the decoupling radius $r_{\rm dec}$ turns out to be insensitive to $M_{\bullet},\alpha,\mu$, and only has a mild dependence on the SMBH accretion rate $\dot{M}_{\bullet}$. From Eq. (14), it is straightforward to see that the sBH is nearly aligned with the disk ($\vec{L}/L\approx\vec{n}_{\rm disk}$) for $r>r_{\rm dec}$ and starts to precess around the SMBH spin axis at $r<r_{\rm dec}$ (as illustrated in FIG. 2). The transition between the two regimes is very sharp due to the sharp $r$ dependence of the two rates $\omega_{\rm LT}$ and $\omega_{\rm DI}$.
@@
respectively. Therefore, wet EMRIs with a warped disk are a natural mechanism to generate type II QPEs. If type II QPEs are confirmed in future observations, their existence provides indirect evidence for warped disks and the chaotic accretion mode in AGNs.
As the EMRI enters the LISA sensitivity band with orbital radius $r\sim 10M_{\bullet}$ and orbital frequency $f_{\rm obt}\sim$ mHz, the energy budgets per collision for the two disk models are quite different. In the $\alpha$-disk model, it scales as $\delta E_{\rm sBH}\propto r^{5/2}\propto f_{\rm obt}^{-5/3}$, that is to say, the mHz QPEs will be $\mathcal{O}(10^{3})$ times weaker than their low-frequency counterparts. Although the identification of mHz type II QPEs is difficult in a sea of much stronger AGN variability, the joint GW measurement of the orbits may help identify the detection of mHz type II QPEs. If the accretion disk can be modeled as a $\beta$-disk, the energy budget per collision scales as $\delta E_{\rm sBH}\propto r^{2/5}\propto f_{\rm obt}^{-4/15}$, i.e. the mHz QPEs will be a few times weaker than their low-frequency counterparts. In this case, the identification of mHz QPEs is possible even without the input of GW observations.
@@
The other possible phenomenon to observe is the strong lensing of flares processed by the central SMBH, similar to self-lensing flares from SMBH binaries [36]. In this case, the strong lensing of the EMRI-produced flares is probably more likely to happen because the sBH is much closer to its host. As shown in Ref. [76], the strong lensing probability of a flare (either an EM flare or GWs) emitted at radius $r$ is approximately
@@
Therefore the chance of strong lensing of a single flare is around $\mathcal{O}(1\%)$ level, and the chance will be largely enhanced considering multi-flares are produced as the EMRI orbiting around the SMBH and crossing the accretion disk at different locations. Unlike the case of SMBH binaries with one of them hosting a mini-accretion disk where the background light source is stable and the variations in the light curve mainly come from when the mini disk is lensed by the foreground BH, the light curve of mHz QPEs varies from flare to flare even without lensing. A more detailed study is required to confirm whether the lensing effect can be identified from the QPE light curve.